OpenStudy (anonymous):

5x-9=3x+1 please help not just answer please have steps.

OpenStudy (anonymous):

A. 4 B. 5 C. 8 D. 10

OpenStudy (anonymous):

Add 9 to both sides 5x=3x+1+9 Simplify 3x+1+9 to 3x+10 5x=3x+10 Subtract 3x from both sides 5x−3x=10 Simplify 5x−3x to 2x 2x=10 Divide both sides by 2 x=10/2 Simplify 10/2 to 5 x=5

OpenStudy (anonymous):

am i right?